One number is 4 more than 3 times another. Their product is 20. Find the numbers
Answer by TimothyLamb(4379)

m = 3n + 4
nm = 20
---
nm = 20
n(3n + 4) = 20
3nn + 4n = 20
3nn + 4n - 20 = 0
---
the above quadratic equation is in standard form, with a=3, b=4, and c=-20
---
to solve the quadratic equation, plug this:
3 4 -20
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots of the quadratic are:
2
-3.33333333 = -(3 + 1/3)
---
there are two solutions to the problem:
n = 2
m = 10
and:
n=-3.33333333
m=-5.99999999
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php